# Geometry-informed Channel Statistics Prediction Based upon Uncalibrated Digital Twins thanks: This research is supported by the HORIZON JU-SNS-2022-STREAM-B-01-02 CENTRIC project (Grant Agreement No.101096379) and the Villum Investigator Grant “WATER” from the Velux Foundations, Denmark. Authors Mahmoud Saad Abouamer and Robin J. Williams contributed equally to this work

### v1

@@ III-A Approximating the statistics of small-scale fading via frequency-domain sampling
Figure 2: Frequency based and location based CDF of received power (left) near transmitter and (right) far from the transmitter.
In Sec. II and [11], it was argued the small-scale fading distribution, that results from small changes in location, can be approximated using the wideband frequency-domain channel power distribution.
@@
We verify this claim using the geometry-consistent DT of the scene. For each of the 127 transmitter locations, the small-scale fading distribution is computed by generating a grid of 400 small movements (on the order of a wavelength $\lambda$) in the transmitter’s $x$ and $y$ coordinates. Using the channel impulse response (CIR) generated by the DT, an empirical location-based distribution of the channel power in dB is computed. The empirical frequency-based channel power distribution is obtained by fixing the transmitter’s position and sampling the frequency domain over a large bandwidth (i.e., 8 GHz, as suggested by [11]).
@@
Fig. 2 shows the empirical cumulative distribution functions (CDFs) of received power for both the frequency-based and location-based approaches. The results are shown for two typical transmitter positions (a) near the transmitter (position 55 in Fig. 1) and (b) far from the transmitter (position 111 in Fig. 1). As Fig. 2 shows, the frequency-domain channel power distribution closely approximates the location-based channel power distribution, including the tails of the distributions. For example, when the receiver is far from the transmitter, the 0.01-quantiles of the location-based and frequency-based distributions differ by only 0.02 dB. Thus, frequency-domain samples, which are can be obtained by wideband channel sounding, offers a practical method to approximate small-scale fading distributions that are otherwise challenging to capture directly [11].
@@
IV Site-specific Channel Statistics Prediction
For the scenario considered in Fig. 1, let $p(\mathbf{x})$ indicate the instantaneous received power associated a transmitter located at position $\mathbf{x}$. Due to small scale fading, $p(\mathbf{x})$ is modeled as a random variable. As demonstrated in [11], reliable rate selection can be achieved using accurate estimates of the statistics of $p(\mathbf{x})$. Particularly, The goal is to estimate the log $\epsilon$-quantile $q_{\epsilon}(\mathbf{x})$
$\rho_{\epsilon}(\mathbf{x})$ $=\sup\{\rho>0|P(p(\mathbf{x})\leq\rho)<\epsilon)\}$
$q_{\epsilon}(\mathbf{x})$ $=\ln(\rho_{\epsilon}(\mathbf{x}))$ (9)
For each position $\mathbf{x}$, a DT can ideally be used to estimate the CDF of the received power ${\mathbf{\psi}}^{\sf{DT}}(\beta,\mathbf{x})=P(p(\mathbf{x})\leq\beta)$ which can then be directly used to provide an estimate $q^{\sf{direct}}_{\epsilon}(\mathbf{x})$. However, in this work, we consider uncalibrated DTs generated from open-source maps (e.g. OpenStreetMap [7]) and with default material properties. Thus, due to mismatched material properties, changes in the wireless environment and other non-idealities, the direct use of a DT does not necessarily provide an accurate estimate of $q_{\epsilon}(\mathbf{x})$. However, the DT’s received power CDF ${\mathbf{\psi}}^{\sf{DT}}(\beta,\mathbf{x})$ can be leveraged to provide site-specific geometric insights. We propose to exploit the geometric information implicitly conveyed by ${\mathbf{\psi}}^{\sf{DT}}(.)$ and a small number of channel measurements to predict fading statistics across the entire site.
Let $\mathcal{D}=\{\mathbf{x}_{d},q_{\epsilon}(\mathbf{x}_{d})\}_{d=1}^{D}$ be a data set of log-fading power quantiles formed using $D$ measurements. The quantile function $q_{\epsilon}(\mathbf{x})$ is modeled as GP. As demonstrated in [11], GP regression can be used to predict fading statistics by spatially interpolating channel measurements using a predictive distribution
$q_{\epsilon}^{\sf{spatial}}(\mathbf{x}|\mathcal{D})\sim\mathcal{N}(\mu(\mathbf{x}|\mathcal{D}),\sigma^{2}(\mathbf{x}|\mathcal{D})).$ (10)
In this work, instead of relying on spatial features alone, we employ GP regression to predict fading statistics using geometric features derived from the DT. In particular, by processing the DT’s distribution of received power $\mathbf{\psi}^{\sf{DT}}(\beta,\mathbf{x})$, we employ a predictive distribution
$q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})$ $\sim\mathcal{N}(\mu_{\sf{DT}}(\mathbf{y}|\mathcal{D}),\sigma_{\sf{DT}}^{2}(\mathbf{y}|\mathcal{D})),$ (11)
Here $\mathbf{f}(\cdot)$ uniformly samples $100$ points from the DT’s CDF of received power $\mathbf{\psi}^{\sf{DT}}(\beta,\mathbf{x})$. With this $\mathbf{y}$ and the use of an anisotropic GP kernel $k(\mathbf{y},\mathbf{y}^{\prime})$, the GP does not only capture spatial similarities but also implicitly encodes geometric information provided by the DT. Particularly, for two receiver positions $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ that result in similar DT outputs $\mathbf{\psi}^{\sf{DT}}(\beta,\mathbf{x}_{1})\approx\mathbf{\psi}^{\sf{DT}}(\beta,\mathbf{x}_{2})$, the GP’s covariance function ensures that the predicted channel statistics at these two locations are close to each other. Thus, the DT’s geometric features are captured via the GP’s covariance function, embedding geometric correlations that allows efficient prediction of channel statistics using a limited number of measurements.
In Section V, using a measurement campaign conducted in the site shown in Fig. 1, we corroborate the effectiveness of the proposed $q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})$ in predicting channel statistics with high accuracy. Moreover, this proposed prediction framework is shown to provide significant improvements in prediction errors compared to the following benchmarks:
V Numerical results
V-A Channel Statistics Predictions
(a) Prediction errors by (uncalibrated) DT
(b) Prediction errors by baseline GP
(c) Prediction errors by proposed GP with DT features
@@ V-A Channel Statistics Predictions
Figure 4: Absolute Error in predicting $\epsilon$-quantile of fading using DT, GP baseline and Proposed GP.
@@
In this section, we consider $\epsilon=1\%$ and the measurement campaign data associated with the setup in Fig 1 which is split into a (small) training set and an evaluation set. Here, the (training) data set is given $\mathcal{D}=\{\mathbf{x}_{d},q_{\epsilon}(\mathbf{x}_{d})\}_{d=1}^{D}$ with $D=30$. The goal is to predict the $\epsilon$-quantile associated with the remaining $127-30$ positions.
@@
We demonstrate the ability of the proposed prediction framework $q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})\sim\mathcal{N}(\mu_{\sf{DT}}(\mathbf{y}|\mathcal{D}),\sigma^{2}_{\sf{DT}}(\mathbf{y}|\mathcal{D}))$ presented in Section IV. Fig. (3) shows $\epsilon$-quantile of fading under the ground-truth measurements obtained through closest point interpolation. Fig. (3) shows $\mu_{\sf{DT}}(\mathbf{y}|\mathcal{D})$ used to predict $\epsilon$-quantile of fading with $\mathcal{D}$ containing $30$ observations (corresponding locations are indicated with white dots). To quantify estimation errors, compared to ground-truth measurements in Fig. (3), Fig. (4) shows the absolute prediction error incurred by $\mu_{\sf{DT}}(\mathbf{y}|\mathcal{D})$. For comparison, Fig. (4) and Fig. (4) show the absolute prediction error incurred by the two benchmarks described in Section IV: (a) directly using DT for prediction $q^{\sf{direct}}_{\epsilon}(\mathbf{x})$ and (b) using spatial features only $q_{\epsilon}^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$. As shown in the figure, compared to the benchmarks, the proposed method significantly improves prediction accuracy across the entire scene by effectively leveraging the geometric consistency of the DT. For instance, the proposed prediction scheme reduces the median error to $1.8$ dB compared to $3.7$ dB and $3.2$ dB associated with $q^{\sf{direct}}_{\epsilon}(\mathbf{x})$ and $q_{\epsilon}^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$, respectively.
@@
V-B Rate selection under reliability constraints
For a position $\mathbf{x}$, the predictive distribution $q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})$ is employed to select a rate $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ that complies with a meta-probability $\tilde{p}_{\epsilon}(\mathbf{x})$ reliability constraint [9]. In particular, for a confidence parameter $\delta>0$, the goal is to select $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ such that the corresponding meta-probability $\tilde{p}^{\sf{DT}}_{\epsilon}(\mathbf{x})\leq\delta$. Following [11], $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ is selected using the predictive distribution $q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})$ given in (11). Here, the rate is selected as the $\delta$-quantile of $q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})$. In particular, for predictive mean $\mu_{\sf{DT}}(\mathbf{y}|\mathcal{D})$ and predictive variance $\sigma_{\sf{DT}}^{2}(\mathbf{y}|\mathcal{D})$,
$R^{\sf{DT}}$ $(\mathbf{x}|\mathcal{D})=$ (13)
@@ V-B Rate selection under reliability constraints
The predictive distribution $q_{\epsilon}^{\sf{spatial}}(\mathbf{x}|\mathcal{D})\sim\mathcal{N}(\mu(\mathbf{x}|\mathcal{D}),\sigma^{2}(\mathbf{x}|\mathcal{D})$ can be similarly used to select a corresponding rate select $R^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$.
To evaluate the performance of $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$, we compute its ratio with respect to the $\epsilon$-outage capacity $R_{\epsilon}(\mathbf{x})$ computed using the (perfect channel knowledge) ground truth fading distribution. This metric is referred to the normalized rate and is given by
$\hat{R}^{\sf{DT}}_{\epsilon}(\mathbf{x})=\frac{R^{\sf{DT}}(\mathbf{x}|\mathcal{D})}{R_{\epsilon}(\mathbf{x})}.$ (14)
For $\epsilon=1\%$ and $\delta=10\%$, Fig.  (5) shows the CDF of $R^{\sf{DT}}_{\epsilon}(\mathbf{x})$ across the $127-30$ (testing) locations. For comparison, Fig.  (5) also shows the normalized rate $\hat{R}^{\sf{spatial}}_{\epsilon}(\mathbf{x})=R^{\sf{spatial}}(\mathbf{x}|\mathcal{D})/R_{\epsilon}(\mathbf{x})$. As shown in Fig. (5), both $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ and $R^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$ meet the meta-probability reliability criteria. However, the proposed $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ provides significant boost in the selected rate compared to $R^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$ and as shown by the shift in normalized rate in Fig. (5). For instance, across different locations, the average normalized rate is approximately doubled from $\bar{\hat{R}}^{\sf{spatial}}_{\epsilon}(\mathbf{x})=0.27$ from to $\bar{\hat{R}}^{\sf{DT}}_{\epsilon}(\mathbf{x})=0.45$. Thus, compared to $q^{\sf{spatial}}_{\epsilon}(\mathbf{x}|\mathcal{D})$, $q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})$ allows selection of higher rates while complying with reliability constraints.
Figure 5: Empirical distribution of $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ and $R^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$ normalized by $\epsilon$-outage capacity $R_{\epsilon}(\mathbf{x})$ for $\epsilon=1\%$ and $\delta=10\%$.
@@ VI Conclusions
In this paper we have shown that by incorporating raytraced channel information from an inaccurate DT, one can improve the accuracy of channel statistics prediction. The proposed system is capable of almost doubling the provided data rate while maintaining the same reliability, when compared with the baseline. This increase does not come at the expense of increased overhead and lengthy DT calibration procedures, as the system utilizes an uncalibrated DT based on geometry from OpenStreetMap [7] with all objects, ground included, having material parameters equal to that of a brick wall. The proposed framework successfully utilizes a DT based on freely available geometry data without calibration to directly improve the performance of a communication system.

### v2

@@ IV-B Numerical Validation
Figure 2: Frequency-based and location-based CDF of received power for two TX positions, one near the receiver (position 55 in Fig. 1) and another far from the receiver (position 111).
@@
For each of the 127 transmitter locations shown in Fig. 1, the small-scale fading distribution is computed by employing the DT to generate a grid of 400 small movements (on the order of a wavelength $\lambda$) in the transmitter’s $x$ and $y$ coordinates. Using the channel impulse response (CIR) generated by the DT, an empirical location-based distribution of the channel power in dB is computed. The empirical frequency-based channel power distribution is also obtained using the DT by fixing the transmitter’s position and sampling the frequency domain over a large bandwidth (i.e., 8 GHz, as suggested by [6]).
@@
Fig. 2 shows the empirical cumulative distribution functions (CDFs) of received power for both the frequency-based and location-based approaches. The results are shown for two typical transmitter positions, one near the receiver and another far from the receiver. As Fig. 2 shows, the frequency-domain channel power distribution closely approximates the location-based channel power distribution, including the tails of the distributions. For example, when the receiver is far from the transmitter, the 0.01-quantiles of the location-based and frequency-based distributions differ by only 0.02 dB. Thus, frequency-domain samples, obtained by wideband channel sounding, offer a practical method to approximate small-scale fading distributions that are otherwise challenging to capture directly.
V Experimental results
V-A Experimental Setup and Frequency Measurements
At each of the 127 transmitter locations shown in Fig. 1, wideband channel measurements were performed in the frequency domain. A total of 8 001 equally spaced frequency samples were taken per site, spanning the band from $2$GHz to $10$GHz, resulting in a frequency resolution of $1$MHz. During each measurement, the receiver remained stationary while the transmitter was sequentially positioned at the designated locations. Detailed information on the hardware, and measurement protocols can be found in [6].
@@ V-A Experimental Setup and Frequency Measurements
V-B Channel Statistics Predictions
@@
Figure 4: Absolute Error in predicting $\epsilon$-quantile of fading using (a) (uncalibrated) DT (b) baseline GP and (c) Proposed GP.
@@
In this section, we consider $\epsilon=1\%$ and the measurement campaign consisting of wideband channel measurements for the 127 positions associated with the setup in Fig. 1 which is split into a (small) training set of $30$ positions and an evaluation set of $97$ positions. Here, the (training) data set is given $\mathcal{{D}}=\{\widehat{q}_{\epsilon,d},\mathbf{x}_{d}\}_{d=1}^{D}$ with $D=30$. The goal is to predict the $\epsilon$-quantile associated with the remaining $97$ positions.
@@
We demonstrate the ability of the proposed prediction framework $q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})\sim\mathcal{N}(\mu_{\sf{DT}}(\mathbf{y}|\mathcal{D}),\sigma^{2}_{\sf{DT}}(\mathbf{y}|\mathcal{D}))$ presented in Section III. Fig. 3 shows $\epsilon$-quantile of fading under the ground-truth measurements obtained through closest point interpolation. Fig. 3 shows $\mu_{\sf{DT}}(\mathbf{y}|\mathcal{D})$ used to predict $\epsilon$-quantile of fading with $\mathcal{D}$ containing $30$ observations (corresponding positions are indicated with white dots). To quantify estimation errors, compared to ground-truth measurements in Fig. 3, Fig. 4 shows the absolute prediction error incurred by $\mu_{\sf{DT}}(\mathbf{y}|\mathcal{D})$. For comparison, Fig. 4 and Fig. 4 show the absolute prediction error incurred by the two benchmarks described in Section III: (a) directly using DT for prediction $q^{\sf{direct}}_{\epsilon}(\mathbf{x})$ and (b) using spatial features only $q_{\epsilon}^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$. As shown in the figure, compared to the benchmarks, the proposed method significantly improves prediction accuracy across the entire scene by effectively leveraging the geometric consistency of the DT. For instance, the proposed prediction scheme reduces the median error to $1.8$ dB compared to $3.7$ dB and $3.2$ dB associated with $q^{\sf{direct}}_{\epsilon}(\mathbf{x})$ and $q_{\epsilon}^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$, respectively.
@@
V-C Rate selection under reliability constraints
For a position $\mathbf{x}$, the predictive distribution $q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})$ is employed to select a rate $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ that complies with a meta-probability $\tilde{p}_{\epsilon}(\mathbf{x})$ reliability constraint. In particular, for a confidence parameter $\delta>0$, the goal is to select $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ such that the corresponding meta-probability $\tilde{p}^{\sf{DT}}_{\epsilon}(\mathbf{x})\leq\delta$. Following [6], $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ is selected using the predictive distribution $q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})$ given in (III-B). Here, the rate is selected as the $\delta$-quantile of $q^{\sf{DT}}_{\epsilon}(\mathbf{x}|\mathcal{D})$. In particular, for predictive mean $\mu_{\sf{DT}}(\mathbf{y}|\mathcal{D})$ and predictive variance $\sigma_{\sf{DT}}^{2}(\mathbf{y}|\mathcal{D})$,
@@ V-C Rate selection under reliability constraints
The predictive distribution $q_{\epsilon}^{\sf{spatial}}(\mathbf{x}|\mathcal{D})\sim\mathcal{N}(\mu(\mathbf{x}|\mathcal{D}),\sigma^{2}(\mathbf{x}|\mathcal{D})$ can be similarly used to select a corresponding rate select $R^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$. To evaluate the performance of $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$, we compute its ratio with respect to the $\epsilon$-outage capacity $R_{\epsilon}(\mathbf{x})$ computed using the (perfect channel knowledge) ground truth fading distribution. This metric is referred to the normalized rate and is given by $\hat{R}^{\sf{DT}}_{\epsilon}(\mathbf{x})=\frac{R^{\sf{DT}}(\mathbf{x}|\mathcal{D})}{R_{\epsilon}(\mathbf{x})}$.
For $\epsilon=1\%$ and $\delta=10\%$, Fig. 5 shows the CDF of $\hat{R}^{\sf{DT}}_{\epsilon}(\mathbf{x})$ across the $97$ testing positions. For comparison, the figure also shows the CDF associated with the the normalized rate $\hat{R}^{\sf{spatial}}_{\epsilon}(\mathbf{x})=R^{\sf{spatial}}(\mathbf{x}|\mathcal{D})/R_{\epsilon}(\mathbf{x})$. As shown in the figure, both $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ and $R^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$ satisfy the meta-probability reliability criteria. This follows as the proportion of normalized rates exceeding $1$, indicating violations of the outage capacity, is constrained within the $\delta=10\%$ target. However, the proposed $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ offers a significant boost in the selected rate compared to $R^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$, as evidenced by the overall increase in normalized rates. Specifically, across different positions, the average normalized rate nearly doubles, improving from $\bar{\hat{R}}^{\sf{spatial}}_{\epsilon}(\mathbf{x})=0.27$ to $\bar{\hat{R}}^{\sf{DT}}_{\epsilon}(\mathbf{x})=0.45$. Consequently, despite limited knowledge of the channel or scene properties, the proposed scheme not only upholds the $\delta=10\%$ meta-probability guarantee but also significantly reduces the gap to the ideal benchmark that assumes perfect channel knowledge.
Figure 5: Empirical distribution of $R^{\sf{DT}}(\mathbf{x}|\mathcal{D})$ and $R^{\sf{spatial}}(\mathbf{x}|\mathcal{D})$ normalized by $R_{\epsilon}(\mathbf{x})$ for $\epsilon=1\%$ and $\delta=10\%$.
@@ VI Conclusions
In this paper we have shown that by incorporating ray-traced channel information from an inaccurate DT, one can improve the accuracy of channel statistics prediction. The proposed system is capable of almost doubling the provided data rate while maintaining the same reliability, when compared with the baseline. This increase does not come at the expense of increased overhead and lengthy DT calibration procedures, as the system utilizes an uncalibrated DT based on geometry from OpenStreetMap [5] with all objects, ground included, having material parameters equal to that of a brick wall. The proposed framework successfully utilizes a DT based on freely available geometry data without calibration to directly improve the performance of a communication system.